# Dress-1-to-3: Single Image to Simulation-Ready 3D Outfit with Diffusion Prior and Differentiable Physics

### v1

@@ Abstract.
Recent advances in large models have significantly advanced image-to-3D reconstruction. However, the generated models are often fused into a single piece, limiting their applicability in downstream tasks. This paper focuses on 3D garment generation, a key area for applications like virtual try-on with dynamic garment animations, which require garments to be separable and simulation-ready. We introduce Dress-1-to-3, a novel pipeline that reconstructs physics-plausible, simulation-ready separated garments with sewing patterns and humans from an in-the-wild image. Starting with the image, our approach combines a pre-trained image-to-sewing pattern generation model for creating coarse sewing patterns with a pre-trained multi-view diffusion model to produce multi-view images. The sewing pattern is further refined using a differentiable garment simulator based on the generated multi-view images. Versatile experiments demonstrate that our optimization approach substantially enhances the geometric alignment of the reconstructed 3D garments and humans with the input image. Furthermore, by integrating a texture generation module and a human motion generation module, we produce customized physics-plausible and realistic dynamic garment demonstrations [^1] Project page: https://dress-1-to-3.github.io/.
@@ 1. Introduction
In recent years, significant advancements in image-to-3D asset reconstruction have been driven by the development of powerful image and video generation models. Among these, multiview diffusion models [Chen et al., 2024c; Liu et al., 2023a; Gao et al., 2024] have emerged as a promising approach, effectively leveraging multiview images as intermediate representations to capture 3D information. When fine-tuned on human datasets, these models generalize well to avatar reconstructions from in-the-wild images [Li et al., 2024e; He et al., 2024a]. However, the generated results are often fused into a single piece, making them unsuitable for downstream tasks such as garment animation and interaction.
@@ 2.1. Multi-view Diffusion
Owing to their powerful predictive ability, Diffusion Probabilistic Models [Ho et al., 2020] have been applied to image [Nichol et al., 2021; Zhang et al., 2023; Dhariwal and Nichol, 2021; Ruiz et al., 2023; Saharia et al., 2022], video [Chen et al., 2024d; Ho et al., 2022], and 3D shape synthesis tasks [Long et al., 2023; Yu et al., 2023; Tang et al., 2023], etc. However, applying image diffusion models to generate multi-view images separately poses significant challenges in maintaining consistency across different views. To address multi-view inconsistency, multi-view attentions and camera pose controls are adopted to fine-tune pre-trained image diffusion models, enabling the simultaneous synthesis of multi-view images [Shi et al., 2023b; Wang and Shi, 2023; Xu et al., 2023b; Yang et al., 2024; Shi et al., 2023a; Long et al., 2024], though these methods might result in compromised geometric consistency due to the lack of inherent 3D biases. To ensure both global semantic consistency and detailed local alignment in multi-view diffusion models, 3D-adapters [Chen et al., 2024a] propose a plug-in module designed to infuse 3D geometry awareness. Nevertheless, the generated images by these models are sparse views. To address this issue, CAT3D [Gao et al., 2024] introduces an efficient parallel sampling strategy to generate a large set of camera poses, and MVDiffusion++ [Tang et al., 2025] adopts a pose-free architecture and a view dropout strategy to reduce computational costs, generating dense, high-resolution images.
Generating consistent images from multi-view diffusions offers guidance for further 3D shape reconstruction [Gao et al., 2024]. PSHuman [Li et al., 2024e] integrates a body-face cross-scale diffusion with an SMPL-X conditioned multi-view diffusion for clothed human reconstruction with high-quality face details. Recent work, MagicMan [He et al., 2024a], utilizes a hybrid human-specific multi-view diffusion model with 3D SMPL-X-based body priors and 2D diffusion priors to consistently generate dense multi-view RGB images and normal maps, supporting high-quality human mesh reconstruction. Different from these works, we exploited multi-view diffusions to generate multi-view normals and RGB images as guidance to optimize sewing patterns and stitches instead of human meshes.
@@ 2.2. Garment Reconstruction
Previous work focusing on clothed human reconstruction [Xiu et al., 2022, 2023] typically generates garments fused with digital human models, limiting them to basic skinning-based animations and requires extra segmentation and editing to separate the garments from the human body. In contrast, our approach focuses on reconstructing separately wearable, simulation-ready garments and human models. Other closely related works include Li et al. [2024a], which also generates simulation-ready clothes, but at the cost of creating clothing templates by artists and precise point clouds by scanners. NeuralTailor [Korosteleva and Lee, 2022] exploits point-level attention for pattern shape and stitching information regression, enabling the reconstruction of garment meshes from point clouds. In contrast, our paper focuses on reconstructing non-watertight garments and humans separately from a single image without additional inputs.
@@
Instead of directly reconstructing garment meshes, some works [Liu et al., 2023b; He et al., 2024b; Korosteleva and Sorkine-Hornung, 2023] treat sewing patterns as intermediate representations to generate garments by stitching them together. Recent work, GarmentRecovery [Li et al., 2024b], introduces implicit sewing patterns (ISP) to provide shape priors integrated with deformation priors for further garment recovery, though it builds specialized models for each individual garment or garment type. Both SewFormer [Liu et al., 2023b] and PanelFormer [Chen et al., 2024b] utilize Transformers to predict sewing patterns and stitches. However, their garment results lack physical material parameters. Therefore, they fail to reconstruct diverse shapes for garments with different physical materials. In addition, these feed-forward methods require large amounts of garment data for training, failing to synthesize garments that fall outside the distribution of the training data. Our work aims to generate diverse, image-aligned, simulation-ready garments with high-quality details from in-the-wild images by optimizing sewing patterns and stitches with physical parameters via differentiable simulations.
@@ 2.3. Differentiable Simulation
Differentiable simulation has seen widespread application in recent research, particularly for system identification and the inference of material parameters from both synthetic [Li et al., 2023a, 2024d] and real-world [Huang et al., 2024; Si et al., 2024] observations. The scope of exploration spans various domains, including fluid dynamics and control [McNamara et al., 2004; Schenck and Fox, 2018; Li et al., 2023b, 2024d], rigid-body dynamics [Freeman et al., 2021; Strecke and Stueckler, 2021; Xu et al., 2023a], articulated systems [Geilinger et al., 2020; Qiao et al., 2021; Xu et al., 2021], soft-body dynamics [Hahn et al., 2019; Hu et al., 2019b; Du et al., 2021; Jatavallabhula et al., 2021; Huang et al., 2024], cloth [Li et al., 2022; Stuyck and Chen, 2023; Li et al., 2024a], inelasticity [Huang et al., 2021; Li et al., 2023a], inflatable structures [Panetta et al., 2021], and Voronoi diagrams [Numerow et al., 2024].
Cloth-based applications, whether for static optimization or dynamic simulation, frequently involve extensive frictional contact. Consequently, many works focus on robust methods for handling dry frictional contact in differentiable simulations. Bartle et al. [2016] proposes a physics-driven pattern adjustment for garment editing using fixed-point optimization, which does not account for gradients. Liang et al. [2019] is the first to introduce a fully functional differentiable cloth simulator with frictional contact and self-collision, formulating a quadratic programming problem. Jatavallabhula et al. [2021] employs a penalty-based frictional contact model, while Du et al. [2021] and Li et al. [2022] leverage the adjoint method for Projective Dynamics [Bouaziz et al., 2014] with friction. Building on Position-Based Dynamics [Müller et al., 2007; Macklin et al., 2016], Stuyck and Chen [2023] and Li et al. [2024a] introduce differentiable formulations for compliant constraint dynamics, and Huang et al. [2024] presents an adjoint-based framework for differentiable Incremental Potential Contact (IPC) [Li et al., 2020, 2021].
@@ 5.1.1. Patch Symmetrization
Here, $\bm{R}_{S}=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$ represents the flip matrix, assuming the symmetry axis is vertical. This optimization involves solving a fixed-coefficient, positive definite linear system, which ensures differentiability. The validated edge control points $\{\hat{\bm{K}}^{e}\}$ are computed analytically by symmetrizing their relative coordinates. The symmetrization constraints are illustrated in the inset figure. Throughout this paper, we omit the hat notation for validated vertices and control points, as all computations are based on the symmetrized patches. However, it is important to note that the underlying garment optimization variables remain the original, non-symmetry-enforced geometry parameters.
@@ 6.4. Post-Iteration Processing
(30) $\mathcal{L}^{\text{Flip}}(\mathcal{P},\mathcal{K})=\frac{1}{|F|}\sum_{f}(\epsilon-\min\{A_{f}(\bm{X}),\epsilon\})+\lambda^{\text{Flip}}\|\bm{X}-\bm{X}^{n+1}\|_{1},$
where $|F|$ is the number of faces $A_{f}$ is the signed area of triangle $f$ and $\bm{X}^{n+1}$ is the discretized garment vertices after the parameter update at iteration $n$. We optimize the above loss only if no triangles are close to flipping.
@@ 7.2. Showcase under Human Motions
The reconstructed simulation-ready garment and human model can be used to generate realistic dynamic human motion in clothing using the robust CIPC physics-based simulator. However, IPC-based simulators require the initial configuration of the human model to be penetration-free. As IPC-based simulators produce intersection-free results, self-penetrations of the human model during the given motion can cause solution failures when the human model interacts with garments. To address this issue, we replace the human model during motion with the nearest intersection-free human model by solving Injective Deformation Processing (IDP) [Fang et al., 2021] problems. In solving these IDP problems, we follow the method in Li et al. [2024c] where the authors use PBD simulations to resolve collisions between garments and human models. Instead, we apply extended Position-Based Dynamics (XPBD) [Macklin et al., 2016] simulations to resolve self-penetrations in the human model by repelling colliding vertices and faces, while preserving natural deformations.
@@ 8. Implementation
We implement CIPC simulation using NVIDIA Warp [Macklin, 2022] to utilize the Auto-Diff feature. The simulator is wraped in a customized autograd.Function to be integrated to the global computational graph.
@@
The pre-optimization steps requires about 10 minutes. The garment optimization process can be finished within 2 hours on a single RTX 3090 with 24GB device memory.
@@ 9.3. Textured Garment Reconstruction and Simulation
To evaluate the generative capability of our method, we perform extensive tests on a variety of images from different sources. We begin by utilizing real-world images from the DeepFashion2 [Ge et al., 2019] dataset, which comprises in-the-wild clothing images sourced from the internet, including those from commercial shops and customers. These images exhibit varying quality and are captured under arbitrary camera and scene settings, reflecting casual user captures. In addition to real-world images, we evaluate our model using synthetic fashion images from the DeepFashion-Multimodal [Jiang et al., 2022] dataset. This dataset features high-resolution clothing images of various types, generated by a diffusion-based transformer. To further extend the generative capability of our method with text prompts, we employ FLUX [Labs, 2023] to generate input images using custom textual descriptions of clothing on a model. For instance, prompts such as "a female model wearing a blazer and pants" are used. To enhance the diversity of the generated results, we randomly incorporate detailed descriptions, including the shape and color of the clothing, as well as the pose and appearance of the model.
@@ 9.4. Ablation Study
Laplacian loss $\mathcal{L}_{\text{Smooth}}$ is applied to smooth out noise and irregular wrinkles in the reconstructed garment mesh. This loss minimizes high-frequency artifacts, enabling a cleaner and more aesthetically pleasing surface. The weight of $\mathcal{L}_{\text{Smooth}}$ is a tunable parameter, allowing users to control the degree of smoothness based on their preferences. As shown in Figure 8, a higher weight results in smoother results but may slightly reduce detail, whereas a lower weight preserves intricate wrinkles but may retain noise.
@@
The boundary corner regularizer, $\mathcal{L}_{\text{BC}}$, mitigates the occurrence of sharp angles in the reconstructed sewing patterns. Sharp or acute angles can lead to practical difficulties during garment fabrication, as they introduce challenges in stitching and material handling. As demonstrated in Figure 8, models trained without $\mathcal{L}_{\text{BC}}$ often generate sewing patterns with acute or impractical corner geometries, whereas incorporating this regularizer results in smoother, more fabrication-friendly boundaries.
@@ 10. Conclusion
In this paper, we present a garment reconstruction pipeline, Dress-1-to-3, which takes a single-view image as input and reconstructs a posed human wearing textured garments, with both the human pose and garment shapes closely aligned with the input image. During optimization, we refine the sewing pattern shapes and physical material parameters by leveraging a differentiable CIPC simulator with accurate frictional contact. The resulting garment assets are simulation-ready and can be seamlessly integrated into a physics-based simulator.

### v2

@@ Abstract.
Recent advances in large models have significantly advanced image-to-3D reconstruction. However, the generated models are often fused into a single piece, limiting their applicability in downstream tasks. This paper focuses on 3D garment generation, a key area for applications like virtual try-on with dynamic garment animations, which require garments to be separable and simulation-ready. We introduce Dress-1-to-3, a novel pipeline that reconstructs physics-plausible, simulation-ready separated garments with sewing patterns and humans from an in-the-wild image. Starting with the image, our approach combines a pre-trained image-to-sewing pattern generation model for creating coarse sewing patterns with a pre-trained multi-view diffusion model to produce multi-view images. The sewing pattern is further refined using a differentiable garment simulator based on the generated multi-view images. Versatile experiments demonstrate that our optimization approach substantially enhances the geometric alignment of the reconstructed 3D garments and humans with the input image. Furthermore, by integrating a texture generation module and a human motion generation module, we produce customized physics-plausible and realistic dynamic garment demonstrations. Our project page is https://dress-1-to-3.github.io/.
@@ 1. Introduction
In recent years, significant advancements in image-to-3D asset reconstruction have been driven by the development of powerful image and video generation models. Among these, multiview diffusion models [Chen et al., 2024c; Liu et al., 2023a; Gao et al., 2024] have emerged as a promising approach, effectively leveraging multiview images as intermediate representations to capture 3D information. When fine-tuned on human datasets, these models generalize well to avatar reconstructions from in-the-wild images [Li et al., 2024d; He et al., 2024a]. However, the generated results are often fused into a single piece, making them unsuitable for downstream tasks such as garment animation and interaction.
@@ 2.1. Multi-view Diffusion
Owing to their powerful predictive ability, Diffusion Probabilistic Models [Ho et al., 2020] have been applied to image [Nichol et al., 2021; Zhang et al., 2023; Dhariwal and Nichol, 2021; Ruiz et al., 2023; Saharia et al., 2022], video [Chen et al., 2024d; Ho et al., 2022], and 3D shape synthesis tasks [Long et al., 2024; Yu et al., 2024b; Tang et al., 2024], etc. However, applying image diffusion models to generate multi-view images separately poses significant challenges in maintaining consistency across different views. To address multi-view inconsistency, multi-view attentions and camera pose controls are adopted to fine-tune pre-trained image diffusion models, enabling the simultaneous synthesis of multi-view images [Shi et al., 2024; Wang and Shi, 2023; Xu et al., 2024; Yang et al., 2024; Shi et al., 2023; Long et al., 2024], though these methods might result in compromised geometric consistency due to the lack of inherent 3D biases. To ensure both global semantic consistency and detailed local alignment in multi-view diffusion models, 3D-adapters [Chen et al., 2024a] propose a plug-in module designed to infuse 3D geometry awareness. Nevertheless, the generated images by these models are sparse views. To address this issue, CAT3D [Gao et al., 2024] introduces an efficient parallel sampling strategy to generate a large set of camera poses, and MVDiffusion++ [Tang et al., 2025] adopts a pose-free architecture and a view dropout strategy to reduce computational costs, generating dense, high-resolution images.
Generating consistent images from multi-view diffusions offers guidance for further 3D shape reconstruction [Gao et al., 2024]. PSHuman [Li et al., 2024d] integrates a body-face cross-scale diffusion with an SMPL-X conditioned multi-view diffusion for clothed human reconstruction with high-quality face details. Recent work, MagicMan [He et al., 2024a], utilizes a hybrid human-specific multi-view diffusion model with 3D SMPL-X-based body priors and 2D diffusion priors to consistently generate dense multi-view RGB images and normal maps, supporting high-quality human mesh reconstruction. Different from these works, we exploited multi-view diffusions to generate multi-view normals and RGB images as guidance to optimize sewing patterns and stitches instead of human meshes.
@@ 2.2. Garment Reconstruction
Previous work focusing on clothed human reconstruction [Xiu et al., 2022, 2023] typically generates garments fused with digital human models, limiting them to basic skinning-based animations and requiring extra segmentation and editing to separate the garments from the human body. In contrast, our approach focuses on reconstructing separately wearable, simulation-ready garments and human models. Other closely related works include Li et al. [2024a]; Yu et al. [2024a], which also generates simulation-ready clothes via differentiable simulation, but at the cost of creating clothing templates by artists, precise point clouds by scanners or 3D shapes of garments. NeuralTailor [Korosteleva and Lee, 2022] utilizes point-level attention for pattern shape and stitching information regression, enabling the reconstruction of garment meshes from point clouds. In contrast, our paper focuses on reconstructing non-watertight garments and humans separately from a single image without additional inputs.
@@
Instead of directly reconstructing garment meshes, some works [Liu et al., 2023b; He et al., 2024b; Korosteleva and Sorkine-Hornung, 2023] treat sewing patterns as intermediate representations to generate garments by stitching them together. Recent work, GarmentRecovery [Li et al., 2024b], introduces implicit sewing patterns (ISP) to provide shape priors integrated with deformation priors for further garment recovery, though it builds specialized models for each individual garment or garment type. Both SewFormer [Liu et al., 2023b], and PanelFormer [Chen et al., 2024b] utilize Transformers to predict sewing patterns and stitches. Concurrent work, SewingLDM, AIpparel, Design2GarmentCode, and ChatGarment [Liu et al., 2024; Nakayama et al., 2024; Zhou et al., 2025; Bian et al., 2025], exploits multimodal models to synthesize sewing patterns. However, their garment results lack physical material parameters. Therefore, they fail to reconstruct diverse shapes for garments with different physical materials. While Wang et al. [2018] and Yang et al. [2018] leverage a shared latent space and joint material-pose optimization to generate 3D garments and 2D sewing patterns, their approaches rely heavily on large-scale datasets on garment templates and human-body models, limiting their ability to generalize to out-of-distribution garments and body shapes. Our work aims to generate diverse, image-aligned, simulation-ready garments with high-quality details from in-the-wild images by optimizing sewing patterns and stitches with physical parameters via differentiable simulations.
@@ 2.3. Differentiable Simulation
Differentiable simulation has seen widespread application in recent research, particularly for system identification and the inference of material parameters from both synthetic [Li et al., 2023a, 2024c] and real-world [Huang et al., 2024; Si et al., 2024] observations. The scope of exploration spans various domains, including fluid dynamics and control [McNamara et al., 2004; Schenck and Fox, 2018; Li et al., 2023b, 2024c], rigid-body dynamics [Freeman et al., 2021; Strecke and Stueckler, 2021; Xu et al., 2023], articulated systems [Geilinger et al., 2020; Qiao et al., 2021; Xu et al., 2021], soft-body dynamics [Hahn et al., 2019; Hu et al., 2019b; Du et al., 2021; Jatavallabhula et al., 2021; Huang et al., 2024], cloth [Li et al., 2022; Stuyck and Chen, 2023; Li et al., 2024a], inelasticity [Huang et al., 2021; Li et al., 2023a], inflatable structures [Panetta et al., 2021], and Voronoi diagrams [Numerow et al., 2024].
Cloth-based applications, whether for static optimization or dynamic simulation [Santesteban et al., 2022; Grigorev et al., 2023], frequently involve extensive frictional contact. Consequently, many works focus on robust methods for handling dry frictional contact in differentiable simulations. Bartle et al. [2016] proposes a physics-driven pattern adjustment for garment editing using fixed-point optimization, which does not account for gradients. Liang et al. [2019] is the first to introduce a fully functional differentiable cloth simulator with frictional contact and self-collision, formulating a quadratic programming problem. Jatavallabhula et al. [2021] employs a penalty-based frictional contact model, while Du et al. [2021] and Li et al. [2022] leverage the adjoint method for Projective Dynamics [Bouaziz et al., 2014] with friction. Building on Position-Based Dynamics [Müller et al., 2007; Macklin et al., 2016], Stuyck and Chen [2023] and Li et al. [2024a] introduce differentiable formulations for compliant constraint dynamics, and Huang et al. [2024] presents an adjoint-based framework for differentiable Incremental Potential Contact (IPC) [Li et al., 2020, 2021].
@@ 5.1.1. Patch Symmetrization
Here, $\bm{R}_{S}=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$ represents the flip matrix, assuming the symmetry axis is vertical. This optimization involves solving a fixed-coefficient, positive definite linear system, which ensures differentiability. The validated edge control points $\{\hat{\bm{K}}^{e}\}$ are computed analytically by symmetrizing their relative coordinates. The symmetrization constraints are illustrated in the inset figure. Throughout this paper, we omit the hat notation for validated vertices and control points, as all computations are based on the symmetrized patches. However, it is important to note that the underlying garment optimization variables retain the original, non-symmetry-enforced geometry parameters.
@@ 6.4. Post-Iteration Processing
(30) $\mathcal{L}_{\text{Flip}}(\mathcal{P},\mathcal{K})=\frac{1}{|F|}\sum_{f}(\epsilon-\min\{A_{f}(\bm{X}),\epsilon\})+\lambda^{\text{Flip}}\|\bm{X}-\bm{X}^{n+1}\|_{1},$
where $|F|$ is the number of faces $A_{f}$ is the signed area of triangle $f$ and $\bm{X}^{n+1}$ is the discretized garment vertices after the parameter update at iteration $n$. We optimize the above loss only if triangles are close to flipping.
@@ 7.2. Showcase under Human Motions
The reconstructed simulation-ready garment and human model can be used to generate realistic dynamic human motion in clothing using the robust CIPC physics-based simulator. However, IPC-based simulators require the initial configuration of the human model to be penetration-free. As IPC-based simulators produce intersection-free results, self-penetrations of the human model during the given motion can cause solution failures when the human model interacts with garments. To address this issue, we replace the human model during motion with the nearest intersection-free human model by solving Injective Deformation Processing (IDP) [Fang et al., 2021] problems. In solving these IDP problems, we follow the method in Li et al. [2025] where the authors use PBD simulations to resolve collisions between garments and human models. Instead, we apply extended Position-Based Dynamics (XPBD) [Macklin et al., 2016] simulations to resolve self-penetrations in the human model by repelling colliding vertices and faces, while preserving natural deformations.
@@ 8. Implementation
We implement CIPC simulation using NVIDIA Warp [Macklin, 2022] to utilize the Auto-Diff feature. The simulator is wrapped in a customized autograd.Function to be integrated to the global computational graph.
@@
The pre-optimization steps require about 10 minutes. The garment optimization process can be finished within 2 hours on a single RTX 3090 with 24GB device memory.
@@ 9.3. Textured Garment Reconstruction and Simulation
To evaluate the generative capability of our method, we perform extensive tests on a variety of images from different sources including 4D-Dress [Wang et al., 2024], CloSe [Antić et al., 2024] and DeepFashion2 [Ge et al., 2019]. These images exhibit diverse quality levels and human poses, highlighting the robustness of our method. To further extend the generative capability of our method with text prompts, we employ FLUX [Labs, 2023] to generate input images using custom textual descriptions of clothing on a model. For instance, prompts such as "a female model wearing a blazer and pants" are used. To enhance the diversity of the generated results, we randomly incorporate detailed descriptions, including the shape and color of the clothing, as well as the pose and appearance of the model.
@@ 9.4. Ablation Study
Laplacian loss $\mathcal{L}_{\text{Lap}}$ is applied to smooth out noise and irregular wrinkles in the reconstructed garment mesh. This loss minimizes high-frequency artifacts, enabling a cleaner and more aesthetically pleasing surface. The weight of $\mathcal{L}_{\text{Lap}}$ is a tunable parameter, allowing users to control the degree of smoothness based on their preferences. As shown in Figure 8, a higher weight results in smoother results but may slightly reduce detail, whereas a lower weight preserves intricate wrinkles but may retain noise.
@@
The boundary corner regularizer, $\mathcal{L}_{\text{BC}}$, mitigates the occurrence of sharp angles in the reconstructed sewing patterns. Sharp or acute angles can lead to practical difficulties during garment fabrication, as they introduce challenges in stitching and material handling. As demonstrated in Figure 8, optimization results obtained without $\mathcal{L}_{\text{BC}}$ often generate sewing patterns with acute or impractical corner geometries, whereas incorporating this regularizer results in smoother, more fabrication-friendly boundaries.
@@
Figure 9. Comparisons between vertex color renderings and texture renderings.
Vertex Color Reconstruction
Vertex colors are optimized to assist garment optimization. However, due to the limited mesh resolution, the visual appearance synthesized with vertex colors tends to be overly smooth. Additionally, colors from adjacent parts can bleed into part boundaries. Comparisons between vertex color renderings and texture renderings are shown in Figure 9. This necessitates an additional module to generate textures that are not constrained by mesh resolution.
Figure 10. Our method tries to find a static garment fit that approximates a non-static garment snapshot or a fit under grasping forces. The input images are generated by GPT-4o.
Loose Garments
Our method optimizes garments in static fitting states under gravity and body supporting forces. For non-static states, such as a snapshot of flowing, or static states influenced by other external forces such as grasping, our garment optimization process attempts to approximate the non-static states by finding nearby static configurations (as shown in Figure 10). However, these static approximations may not reflect the garment’s true geometry. We acknowledge this as a limitation and leave the extension to dynamic states or broader boundary conditions as future work.
@@ 10. Conclusion
In this paper, we present a garment reconstruction pipeline, Dress-1-to-3, which takes a single-view image as input and reconstructs a posed human wearing textured garments, with both the human pose and garment shapes closely aligned with the input image. During optimization, we refine the sewing pattern shapes and physical material parameters by leveraging a differentiable CIPC simulator with accurate contact. The resulting garment assets are simulation-ready and can be seamlessly integrated into a physics-based simulator.
@@
Furthermore, some optimized sewing patterns may not fully adhere to conventional fashion design principles, as our supervision relies solely on ground-truth renderings of fitted garments. Incorporating regularizers based on design conventions in future work could help produce patterns that are more suitable for manufacturing.
Another limitation lies in the overly smoothed garment surfaces produced by our method. To enhance training robustness, we incorporate regularizers such as seam loss and Laplacian loss; however, these also suppress the formation of natural wrinkles. Another contributing factor to the lack of high-frequency detail is the inconsistency in multi-view normal maps generated by MagicMan, which further smooths geometry in detailed regions. Addressing the reconstruction of high-frequency geometric features remains an avenue for future work.
We also observe a gap between input images and generated textures. Since this is not our primary technical contribution, we use off-the-shelf tools for texture generation. Improving PBR texture generation is left for future work, as it warrants a standalone research effort.
Lastly, although our simulation layer supports differentiable dynamic simulation, we currently use it only for static fittings under gravity and body support forces. Extending the garment optimization to handle dynamic scenarios, such as reconstruction from monocular videos, or interaction-driven deformations like grasping, would be both interesting and practically valuable.
Ethical Concerns
We acknowledge that body shape biases exist in our input image datasets.